# The cone conjecture for primitive symplectic varieties over a field of characteristic zero and an application

### v1

@@ 1. Introduction
    There exists a rational polyhedral cone $\Pi$ which is a fundamental domain for the action of the group of relative pseudoautomorphisms $\mathop{\rm PsAut}\nolimits(X/S)$ on $\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S):=\mathop{\overline{\rm Mov}}\nolimits(X/S)\cap\mathop{\rm Eff}\nolimits(X/S)$ in the following sense:
@@
    There exists a rational polyhedral cone $\Pi^{\prime}$ which is a fundamental domain for the action of the group of relative automorphisms $\mathop{\rm Aut}\nolimits(X/S)$ on $\mathop{\rm Nef}\nolimits^{e}(X/S):=\mathop{\rm Nef}\nolimits(X/S)\cap\mathop{\rm Eff}\nolimits(X/S)$ in the above sense.
@@
Beyond its intrinsic interest, the Kawamata–Morrison cone conjecture has important implications for the birational geometry of varieties with trivial canonical divisor, see [undefy, Theorem 1.5]. Very little is known in the relative setting. We may nevertheless mention the work of Y. Kawamata in dimension three [undefam], as well as the recent works [undefav], [undefau] and [undefaaf], which establish weak versions of the relative cone conjecture for fibrations in surface.
@@
The absolute cone conjecture has been verified in a number of cases, but is far from being established in full generality and remains open in dimension at least three. For example, the case of smooth surfaces (with boundaries) is covered by the work of H. Sterk [undefaas], Y. Namikawa [undefaai] and B. Totaro [undefaav]. The case of abelian varieties has been dealt with entirely in [undefaan]. More recently, thanks to the work of numerous geometers, this conjecture has been proven for an irreducible holomorphic symplectic manifold $X$ when the effective nef cone $\mathop{\rm Nef}\nolimits^{e}(X)$ is replaced by the convex hull of $\mathop{\rm Nef}\nolimits(X)\cap N^{1}(X)_{\mathbb{Q}}$ which is usually denoted by $\mathop{\rm Nef}\nolimits^{+}(X)$, see for example [undef], [undefa], [undefaw] and [undefay]. In general, Boucksom-Zariski divisorial decomposition allows us to show that $\mathop{\rm Nef}\nolimits^{e}(X)\subset\mathop{\rm Nef}\nolimits^{+}(X)$, but the reverse inclusion is more difficult to prove when we do not know whether integral nef isotropic classes (with respect to the Beauville-Bogomolov-Fujiki form) are effective. However, given that the SYZ conjecture for irreducible holomorphic symplectic manifolds [undefaao, Conjecture 4.1] has been proven for the four known families ([undefe, Theorem 1.5], [undefax, Theorem 1.3], [undefaaz, Proposition 3.38]), [undefaz, Corollary 1.1] [undefaae, Theorem 7.2, Corollary 7.3] and [undefaad, Theorem 2.2]) we have in these cases the equality $\mathop{\rm Nef}\nolimits^{e}(X)=\mathop{\rm Nef}\nolimits^{+}(X)$ and consequently, the Kawamata-Morrison conjectures are proven for these families. The relative movable cone conjecture has been proven in [undefai] for a fibration whose very general fibre is a projective irreducible holomorphic symplectic manifold, assuming that the total space is klt and assuming that every nef divisor on the very general fibre is semiample.
@@
The absolute cone conjecture has also been established for primitive symplectic varieties (see Definition 2.1), which are the singular analogues of complex irreducible holomorphic symplectic manifolds, under suitable assumptions, see [undefat, Theorem 1.2]. These varieties are equally interesting given that they form one of the three classes along with complex tori and singular Calabi-Yau varieties that can be used to express, up to a quasi-étale covering, any varieties with log-terminal singularities and trivial first Chern class, see for example [undefu], [undeft], [undefac], [undefaj] and [undefc]. Such varieties may appear when considering certain moduli spaces; for example, if $S$ is a K3 surface or if $A$ is an abelian surface, for a fixed Mukai vector $v$ on $S$ or $A$, the spaces $M_{v}(S,H)$ and $K_{v}(A,H)$ are almost always primitive symplectic varieties with terminal singularities (see for instance [undefaam, Theorem 1.10] and the paragraph following the statement, as well as Section 1 for notations). We can also mention Nikulin orbifolds: starting from an irreducible holomorphic symplectic variety $X$ of type $K3^{[2]}$ with a symplectic involution $\iota\in\mathop{\rm Aut}\nolimits(X)$, the quotient $\hat{X}=X/\iota$ is singular along 28 isolated points and a K3 surface denoted $\Sigma$. The blow-up of $\hat{X}$ along $\Sigma$ is then a primitive symplectic variety called a Nikulin orbifold, see [undefaab, Proposition 3.8]. The first result of this paper is the proof of the absolute nef and movable cone conjectures for projective primitive symplectic varieties over a field of characteristic $0$.
@@
Let $X$ be a projective primitive symplectic variety over a field $F$ of characteristic $0$ such that $b_{2}(X)\geq 5$ and $X_{\overline{F}}$ has $\mathbb{Q}$-factorial and terminal singularities.
@@
    The absolute movable cone conjecture as stated in 1.2, item $(1)$, holds for $X$ and the $\mathop{\rm Bir}\nolimits(X)$-action on $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$. More precisely, there exists a rational polyhedral cone $\Pi$ which is a fundamental domain for the action of $\mathop{\rm Bir}\nolimits(X)$ on $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$.
@@
    The absolute nef cone conjecture as stated in 1.2, item $(2)$, holds for $X$ and the $\mathop{\rm Aut}\nolimits(X)$-action on $\mathop{\rm Nef}\nolimits^{+}(X)$. More precisely, there exists a rational polyhedral cone $\Pi^{\prime}$ which is a fundamental domain for the action of $\mathop{\rm Aut}\nolimits(X)$ on $\mathop{\rm Nef}\nolimits^{+}(X)$.
Theorem A is a generalization to the singular case of [undefaat, Theorem 1.0.5] and its proof follows the same approach that goes back to [undefaas], [undefaw] and [undefm]. This theorem can be used to prove the next result that is a natural generalization of [undefai, Theorem 1.1] to the setting of singular fibres.
@@
    the very general fibre of $f$ is a projective primitive symplectic variety with $\mathbb{Q}$-factorial, terminal singularities and $b_{2}(F)\geq 5$,
@@
The proof of this theorem follows the same strategy as in [undefai]. The issue of the existence of both minimal models and good minimal models for primitive symplectic varieties is an active subject of research. The recent work by C. Onorati and Á.D. Ríos Ortiz [undefaal, Theorem A] provides a proof of the semiampleness of nef divisors on primitive symplectic varieties which are locally trivial deformations of $M_{v}(S,H)$, with $v=m\cdot w$ a non-primitive Mukai, under natural numerical hypothesis on $w$. Similarly, thanks to L. Buelli [undefo, Theorem 5.2.1], the semiampleness of nef divisors is now known on primitive symplectic varieties which are locally trivial deformations of $K_{v}(A,H)$ under similar numerical assumptions. In particular, for fibrations whose very general fibre is a locally trivial deformation of these moduli spaces, our Theorem B applies as soon as termination is known. By [undefas, Theorem 1.3], any effective klt pair $(X,\Delta)$ with $X$ a projective primitive symplectic variety with terminal hyperquotient singularities and $\Delta$ an effective divisor, the pair $(X,\Delta)$ has a minimal model. The termination is therefore known for Nikulin orbifolds [undefas, Corollary 1.4], and G. Nanni proved in [undefaaj] the semiampleness of nef divisors on these primitive symplectic varieties, which provides an example to which our theorem B applies.
During the preparation of this work, Theorem A was stated in [undefx, Theorem 3.5] and its proof very briefly sketched along the lines that we were also naturally and independently following. We hope that the full details provided here as well as the application will be a useful addition to the literature for the interested reader.
@@
In Section 2.1, we give the main definitions concerning primitive symplectic varieties over a field of characteristic 0. We prove some properties of base change with respect to field extensions for these varieties. In Section 2.2, we recall the definition and basic properties of the Picard scheme that we will need to apply Lefschetz principle. This will enable us to equip the torsion-free part of the Picard group of a primitive symplectic variety defined over a field of characteristic zero with a lattice structure. In Section 3.1, we recall the definitions of the various cones with which we will be working and prove some preliminary results for the absolute conjectures. In Section 3.2 and Section 3.3, we prove Theorem A item (1) and Theorem A item (2) respectively. In Section 4.1, we recall some definitions from birational geometry and the Minimal Model Program that will be useful to us. Finally, in Section 4.2, we prove Theorem B.
@@ 2. Definitions and first properties
Throughout this section, we fix a field $F$ of characteristic $0$. If $F\hookrightarrow L$ is a field extension and $X$ is a scheme over $F$, we will denote by $X_{L}$ the $L$-scheme $X\otimes_{F}L:=X\times_{\mathop{\rm Spec}\nolimits F}\mathop{\rm Spec}\nolimits L$.
@@ 2.1. Primitive symplectic varieties, base change and descent
A projective primitive symplectic variety over $F$ is a normal projective variety $X$ over $F$ of dimension $2m=\dim(X)$ such that
@@
Let $F\subset L$ be a field extension and let $X$ be a projective primitive symplectic variety over $F$ of dimension $2m$. Then the base change $X_{L}$ is a projective primitive symplectic variety over $L$ of dimension $2m$.
@@
Next, observe that the regular locus of $X$ commutes with the base change. Indeed, let $p:X_{L}\to X$ be the projection. By construction of the fibre product, we have $p^{-1}(X^{\mathop{\rm reg}\nolimits})\simeq(X^{\mathop{\rm reg}\nolimits})_{L}$. We claim that the open subschemes $(X^{\mathop{\rm reg}\nolimits})_{L}$ and $(X_{L})^{\mathop{\rm reg}\nolimits}$ of $X_{L}$ are equal. We simply need to show that they have the same underlying topological space. The set-theoretically inclusion $(X^{\mathop{\rm reg}\nolimits})_{L}\subset(X_{L})^{\mathop{\rm reg}\nolimits}$ is provided, for example, by [undefab, Proposition 5.1.(4)], and the reverse inclusion $(X_{L})^{\mathop{\rm reg}\nolimits}\subset(X^{\mathop{\rm reg}\nolimits})_{L}$ is always true since $p$ is faithfully flat, see the discussion after [undefab, Remark 18.50]. From now on, we will denote the regular locus of $X_{L}$ by $X_{L}^{\mathop{\rm reg}\nolimits}$.
@@
First of all, let us note that $X^{\prime}$ is (geometrically) normal since $X$ is normal by assumption. Moreover, thanks to [undefad, Corollaire 6.6.5] (note that this result is stated for finite field extensions, but it remains true without this finiteness assumption), we know that $X^{\prime}$ is projective. Now let us fix a resolution of singularities $\pi^{\prime}:Y^{\prime}\to X^{\prime}$ with $Y^{\prime}$ a smooth variety over $F^{\prime}$. Let $Y:=Y^{\prime}_{F}$ and denote by $p_{X}:X\to X^{\prime}$ and $p_{Y}:Y\to Y^{\prime}$ the projections. By base change, we obtain a resolution of singularities $\pi:Y\to X$. The flat base change theorem gives $h^{1}(X^{\prime},\mathcal{O}_{X^{\prime}})=0$ and
$h^{0}\left(X^{\prime},\Omega^{[2]}_{X^{\prime}/F^{\prime}}\right)=1.$
As in the previous proof, the base change induces an isomorphism of algebraic de Rham complexes, so we deduce the existence of a unique, up to multiplication by a nonzero scalar, symplectic 2-form $\sigma^{\prime}$ on $X^{\prime\mathop{\rm reg}\nolimits}$ such that $\sigma:=p_{X}^{*}\sigma^{\prime}$ is a symplectic 2-form on $X^{\mathop{\rm reg}\nolimits}$. The only obstruction to the fact that $X^{\prime}$ is a primitive symplectic primitive variety is that, a priori, $\pi^{\prime*}\sigma^{\prime}$ does not extend to a global $2$-form over $Y^{\prime}$. Set $U^{\prime}:=X^{\prime\mathop{\rm reg}\nolimits}$ and $U:=U^{\prime}_{F}=X^{\mathop{\rm reg}\nolimits}$. Consider the following commutative diagram.
@@
By assumption, there exists $\omega\in\Gamma(Y,\Omega^{2}_{Y/F})$ such that $\omega|_{\pi^{-1}(U)}=\pi^{*}\sigma$. Let $C$ be the cokernel of the restriction map $\Gamma(Y^{\prime},\Omega^{2}_{Y^{\prime}/F^{\prime}})\to\Gamma(\pi^{\prime-1}(U^{\prime}),\Omega^{2}_{Y^{\prime}/F^{\prime}})$ and consider $[\pi^{\prime*}\sigma^{\prime}]\in C$. By the flat base change theorem, the top line of the diagram corresponds to the bottom line tensored by $F$, and under this identification, we have that $\pi^{*}\sigma=\pi^{\prime*}\sigma^{\prime}\otimes 1$ and that $C\otimes_{F^{\prime}}F$ is the cokernel of the restriction map $\Gamma(Y,\Omega^{2}_{Y/F})\to\Gamma(\pi^{-1}(U),\Omega^{2}_{Y/F})$. We can therefore deduce that $[\pi^{\prime*}\sigma^{\prime}]\otimes 1=[\pi^{*}\sigma]=0\in C\otimes_{F^{\prime}}F$, and therefore there exists $\omega^{\prime}\in\Gamma(Y^{\prime},\Omega^{2}_{Y^{\prime}/F^{\prime}})$ such that $\omega^{\prime}|_{\pi^{\prime-1}(U^{\prime})}=\pi^{\prime*}\sigma^{\prime}$ (note that, by injectivity of the restriction $\Gamma(Y,\Omega^{2}_{Y/F})\to\Gamma(\pi^{-1}(U),\Omega^{2}_{Y/F})$, we have $\omega=p_{Y}^{*}\omega^{\prime}$). This shows that $X^{\prime}$ is a primitive symplectic variety over $F^{\prime}$, thus completing the proof. ∎
@@
Proposition 2.5 (Proposition 5.6 - [undefk]).

### v2

@@ 1. Introduction
    There exists a rational polyhedral cone $\Pi$ which is a fundamental domain under the action of the group of relative pseudoautomorphisms $\mathop{\rm PsAut}\nolimits(X/S)$ on $\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S):=\mathop{\overline{\rm Mov}}\nolimits(X/S)\cap\mathop{\rm Eff}\nolimits(X/S)$ in the following sense:
@@
    There exists a rational polyhedral cone $\Pi^{\prime}$ which is a fundamental domain under the action of the group of relative automorphisms $\mathop{\rm Aut}\nolimits(X/S)$ on $\mathop{\rm Nef}\nolimits^{e}(X/S):=\mathop{\rm Nef}\nolimits(X/S)\cap\mathop{\rm Eff}\nolimits(X/S)$ in the above sense.
@@
Beyond its intrinsic interest, the Kawamata–Morrison cone conjecture has important implications for the birational geometry of varieties with trivial canonical divisor, see [undefx, Theorem 1.5]. Very little is known in the relative setting. We may nevertheless mention the work of Y. Kawamata in dimension three [undefal], as well as the recent works [undefau], [undefat] and [undefaaf], which establish weak versions of the relative cone conjecture for fibrations in surface.
@@
The absolute cone conjecture has been verified in a number of cases, but is far from being established in full generality and remains open in dimension at least three. For example, the case of smooth surfaces (with boundaries) is covered by the work of H. Sterk [undefaas], Y. Namikawa [undefaai] and B. Totaro [undefaav]. The case of abelian varieties has been dealt with entirely in [undefaan]. More recently, thanks to the work of numerous geometers, this conjecture has been proven for irreducible holomorphic symplectic manifolds $X$ when the effective nef cone $\mathop{\rm Nef}\nolimits^{e}(X)$ is replaced by the convex hull of $\mathop{\rm Nef}\nolimits(X)\cap N^{1}(X)_{\mathbb{Q}}$ which is usually denoted by $\mathop{\rm Nef}\nolimits^{+}(X)$, see for example [undef], [undefa], [undefaw] and [undefay]. In general, Boucksom-Zariski divisorial decomposition allows us to show that $\mathop{\rm Nef}\nolimits^{e}(X)\subset\mathop{\rm Nef}\nolimits^{+}(X)$, but the reverse inclusion is more difficult to prove when we do not know whether integral nef isotropic classes (with respect to the Beauville-Bogomolov-Fujiki form) are effective. However, given that the SYZ conjecture for irreducible holomorphic symplectic manifolds [undefaao, Conjecture 4.1] has been proven for the four known families ([undefe, Theorem 1.5], [undefax, Theorem 1.3], [undefaaz, Proposition 3.38]), [undefaz, Corollary 1.1] [undefaae, Theorem 7.2, Corollary 7.3] and [undefaad, Theorem 2.2]) we have in these cases the equality $\mathop{\rm Nef}\nolimits^{e}(X)=\mathop{\rm Nef}\nolimits^{+}(X)$ and consequently, the Kawamata-Morrison conjectures are proven for these families. The relative movable cone conjecture has been proven in [undefah] for a fibration whose very general fibre is a projective irreducible holomorphic symplectic manifold, assuming that the total space is klt and assuming that every nef divisor on the very general fibre is semiample.
@@
The absolute cone conjecture has also been established for complex projective primitive symplectic varieties (see Definition 2.1), which are the singular analogues of complex projective irreducible holomorphic symplectic manifolds, under suitable assumptions in [undefas, Theorem 1.2]. These varieties are equally interesting given that they form one of the three classes along with complex tori and singular Calabi-Yau varieties that can be used to express, up to a quasi-étale covering, any varieties with log-terminal singularities and trivial first Chern class, see for example [undeft], [undefs], [undefab], [undefai] and [undefc]. Such varieties may appear when considering certain moduli spaces; for example, if $S$ is a K3 surface or if $A$ is an abelian surface, for a fixed Mukai vector $v$ on $S$ or $A$, the spaces $M_{v}(S,H)$ and $K_{v}(A,H)$ are almost always primitive symplectic varieties with terminal singularities (see for instance [undefaam, Theorem 1.10]). We can also mention Nikulin orbifolds: starting from an irreducible holomorphic symplectic variety $X$ of type $K3^{[2]}$ with a symplectic involution $\iota\in\mathop{\rm Aut}\nolimits(X)$, the quotient $\hat{X}=X/\iota$ is singular along 28 isolated points and a K3 surface denoted $\Sigma$. The blow-up of $\hat{X}$ along $\Sigma$ is then a primitive symplectic variety called a Nikulin orbifold, see [undefaab, Proposition 3.8]. The first result of this paper is the proof of the absolute nef and movable cone conjectures for projective primitive symplectic varieties over a field of characteristic $0$.
@@
Let $X$ be a projective primitive symplectic variety over a field $F$ of characteristic $0$ such that $b_{2}(X)\geq 5$ and $X_{\overline{F}}$ has $\mathbb{Q}$-factorial terminal singularities.
@@
    The absolute movable cone conjecture as stated in 1.2, item $(1)$, holds for $X$ and the $\mathop{\rm Bir}\nolimits(X)$-action on $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$. More precisely, there exists a rational polyhedral cone $\Pi$ which is a fundamental domain under the action of $\mathop{\rm Bir}\nolimits(X)$ on $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$.
@@
    The absolute nef cone conjecture as stated in 1.2, item $(2)$, holds for $X$ and the $\mathop{\rm Aut}\nolimits(X)$-action on $\mathop{\rm Nef}\nolimits^{+}(X)$. More precisely, there exists a rational polyhedral cone $\Pi^{\prime}$ which is a fundamental domain under the action of $\mathop{\rm Aut}\nolimits(X)$ on $\mathop{\rm Nef}\nolimits^{+}(X)$.
Theorem A is a generalization to the singular case of [undefaat, Theorem 1.0.5] and its proof follows the same approach that goes back to [undefaas], [undefaw] and [undefl]. This theorem can be used to prove the next result that is a natural generalization of [undefah, Theorem 1.1] to the setting of singular fibres.
@@
    the very general fibre $X_{s}$ of $f$ is a projective primitive symplectic variety with $b_{2}(X_{s})\geq 5$ and $\mathbb{Q}$-factorial terminal singularities,
@@
The proof of this theorem follows the same strategy as in [undefah]. The issue of the existence of both minimal models and good minimal models for primitive symplectic varieties is an active subject of research. The recent work by C. Onorati and Á.D. Ríos Ortiz [undefaal, Theorem A] provides a proof of the semiampleness of nef divisors on primitive symplectic varieties which are locally trivial deformations of $M_{v}(S,H)$, with $v=m\cdot w$ a non-primitive Mukai, under natural numerical hypothesis on $w$. Similarly, thanks to L. Buelli [undefn, Theorem 5.2.1], the semiampleness of nef divisors is now known on primitive symplectic varieties which are locally trivial deformations of $K_{v}(A,H)$ under similar numerical assumptions. In particular, for fibrations whose very general fibre is a locally trivial deformation of these moduli spaces, our Theorem B applies as soon as termination is known. By [undefar, Theorem 1.3], any effective klt pair $(X,\Delta)$ with $X$ a projective primitive symplectic variety with terminal hyperquotient singularities and $\Delta$ an effective divisor has a minimal model. The termination is therefore known for Nikulin orbifolds [undefar, Corollary 1.4], and G. Nanni proved in [undefaaj] the semiampleness of nef divisors on these primitive symplectic varieties, which provides an example to which our theorem B applies.
During the preparation of this work, Theorem A was stated in [undefw, Theorem 3.5] and its proof very briefly sketched along the lines that we were also naturally and independently following. We hope that the full details provided here as well as the application will be a useful addition to the literature for the interested reader.
@@
In Section 2.1, we give the main definitions concerning primitive symplectic varieties over a field of characteristic 0. We prove some properties of base change with respect to field extensions for these varieties. In Section 2.2, we recall the definition and basic properties of the Picard scheme that we will need to apply the Lefschetz principle. This will enable us to equip the torsion-free part of the Picard group of a primitive symplectic variety defined over a field of characteristic zero with a lattice structure. In Section 3.1, we recall the definitions of the various cones with which we will be working and prove some preliminary results for the absolute conjectures. In Section 3.2 and Section 3.3, we prove Theorem A item (1) and Theorem A item (2) respectively. In Section 4.1, we recall some definitions from birational geometry and the Minimal Model Program that will be useful to us. Finally, in Section 4.2, we prove Theorem B.
@@ 2. Definitions and first properties
Throughout this section, we fix a field $F$ of characteristic $0$. If $F\subset L$ is a field extension and $X$ is a scheme over $F$, we will denote by $X_{L}$ the $L$-scheme $X\otimes_{F}L:=X\times_{\mathop{\rm Spec}\nolimits F}\mathop{\rm Spec}\nolimits L$.
@@ 2.1. Primitive symplectic varieties, base change and descent
A projective primitive symplectic variety over $F$ is a normal projective variety $X$ over $F$ of dimension $2n=\dim(X)$ such that
@@
Let $F\subset L$ be a field extension and let $X$ be a projective primitive symplectic variety over $F$ of dimension $2n$. Then the base change $X_{L}$ is a projective primitive symplectic variety over $L$ of dimension $2n$.
@@
Next, observe that the regular locus of $X$ commutes with the base change. Indeed, let $p:X_{L}\to X$ be the projection. By construction of the fibre product, we have $p^{-1}(X^{\mathop{\rm reg}\nolimits})\simeq(X^{\mathop{\rm reg}\nolimits})_{L}$. We claim that the open subschemes $(X^{\mathop{\rm reg}\nolimits})_{L}$ and $(X_{L})^{\mathop{\rm reg}\nolimits}$ of $X_{L}$ are equal. We simply need to show that they have the same underlying topological space. The set-theoretically inclusion $(X^{\mathop{\rm reg}\nolimits})_{L}\subset(X_{L})^{\mathop{\rm reg}\nolimits}$ is provided, for example, by [undefaa, Proposition 5.1.(4)], and the reverse inclusion $(X_{L})^{\mathop{\rm reg}\nolimits}\subset(X^{\mathop{\rm reg}\nolimits})_{L}$ is always true since $p$ is faithfully flat, see the discussion after [undefaa, Remark 18.50]. From now on, we will denote the regular locus of $X_{L}$ by $X_{L}^{\mathop{\rm reg}\nolimits}$.
@@
First of all, note that $X^{\prime}$ is normal since $X$ is geometrically normal. Moreover, thanks to [undefac, Corollaire 6.6.5] (although this result is stated for finite field extensions, it remains true without this finiteness assumption), we know that $X^{\prime}$ is projective. Now fix a resolution of singularities $\pi^{\prime}:Y^{\prime}\to X^{\prime}$ with $Y^{\prime}$ a smooth variety over $F^{\prime}$. Let $Y:=Y^{\prime}_{F}$, and denote by $p_{X}:X\to X^{\prime}$ and $p_{Y}:Y\to Y^{\prime}$ the projections. By base change, we obtain a resolution of singularities $\pi:Y\to X$ such that $\pi=\pi^{\prime}_{F}$. The flat base change theorem gives $h^{1}(X^{\prime},\mathcal{O}_{X^{\prime}})=0$ and $h^{0}(X^{\prime},\Omega^{[2]}_{X^{\prime}/F^{\prime}})=1$. As in the previous proof, the base change induces an isomorphism of algebraic de Rham complexes, so we deduce the existence of a unique (up to multiplication by a nonzero scalar) symplectic 2-form $\sigma^{\prime}$ on $X^{\prime\mathop{\rm reg}\nolimits}$ such that $\sigma:=p_{X}^{*}\sigma^{\prime}$ is a symplectic 2-form on $X^{\mathop{\rm reg}\nolimits}$. The only obstruction to the fact that $X^{\prime}$ is a primitive symplectic primitive variety is that, a priori, $\pi^{\prime*}\sigma^{\prime}$ does not extend to a global $2$-form over $Y^{\prime}$. Set $U^{\prime}:=X^{\prime\mathop{\rm reg}\nolimits}$ and $U:=U^{\prime}_{F}=X^{\mathop{\rm reg}\nolimits}$. Consider the following commutative diagram.
@@
By assumption, there exists $\omega\in\Gamma(Y,\Omega^{2}_{Y/F})$ such that $\omega|_{\pi^{-1}(U)}=\pi^{*}\sigma$. Let $C$ be the cokernel of the restriction map $\Gamma(Y^{\prime},\Omega^{2}_{Y^{\prime}/F^{\prime}})\to\Gamma(\pi^{\prime-1}(U^{\prime}),\Omega^{2}_{Y^{\prime}/F^{\prime}})$ and consider $[\pi^{\prime*}\sigma^{\prime}]\in C$. By the flat base change theorem, the top line of the diagram corresponds to the bottom line tensored by $F$, and under this identification, we have that $\pi^{*}\sigma=\pi^{\prime*}\sigma^{\prime}\otimes 1$ and that $C\otimes_{F^{\prime}}F$ is the cokernel of the restriction map $\Gamma(Y,\Omega^{2}_{Y/F})\to\Gamma(\pi^{-1}(U),\Omega^{2}_{Y/F})$. We deduce that $[\pi^{\prime*}\sigma^{\prime}]\otimes 1=[\pi^{*}\sigma]=0\in C\otimes_{F^{\prime}}F$, and therefore there exists $\omega^{\prime}\in\Gamma(Y^{\prime},\Omega^{2}_{Y^{\prime}/F^{\prime}})$ such that $\omega^{\prime}|_{\pi^{\prime-1}(U^{\prime})}=\pi^{\prime*}\sigma^{\prime}$ (note that, by injectivity of the restriction $\Gamma(Y,\Omega^{2}_{Y/F})\to\Gamma(\pi^{-1}(U),\Omega^{2}_{Y/F})$, we have $\omega=p_{Y}^{*}\omega^{\prime}$). This shows that $X^{\prime}$ is a primitive symplectic variety over $F^{\prime}$, thus completing the proof. ∎
@@
Proposition 2.5 (Proposition 5.6 - [undefj]).